OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x - 2y = -4 2x - y = 1

OpenStudy (anonymous):

x- 2y = -4 x=2y-4 2(2y-4)-y =1 4y-8-y=1 3y-8=1 3y=9 y=3 x-2y=-4 x-2(3)=-4 x-6=-4 x=2 (x,y) = (2, 3)